OpenStudy (anonymous):

please help, factor 9x^2-9x-10

OpenStudy (anonymous):

please help explain the steps

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\] Where a = 9 b= -9 c= -10

OpenStudy (anonymous):

just plug into the formula to find the values of x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait what is a , isnt it 9?

OpenStudy (anonymous):

yes a=9

OpenStudy (anonymous):

i get